OpenStudy (anonymous):

A group of 12 students have formed a club. They have decided to set up an "executive committee" of four of the members of the club. All four of the members of the committee will have equal power. How many ways could the club select four members for the executive committee? multiply then divide 495 or just multiply 11,880

jimthompson5910 (jim_thompson5910):

Compute 12 C 4 to find your answer. Do you know how to do this?

OpenStudy (anonymous):

yes il either multiply 12 x 11 x10 x9 or divide by 6

jimthompson5910 (jim_thompson5910):

Yes you start by multiplying 12 x 11 x 10 x 9 Then you divide this result by 4! = 4*3*2*1 = 24

OpenStudy (anonymous):

I think this is permutation!

jimthompson5910 (jim_thompson5910):

no, all positions are of equal power

jimthompson5910 (jim_thompson5910):

so order doesn't matter

OpenStudy (anonymous):

Oh, just read it again!

OpenStudy (anonymous):

permutation is only when you multiply?

jimthompson5910 (jim_thompson5910):

yes, if order mattered, then the answer would be 12 x 11 x 10 x 9 multiplied out

jimthompson5910 (jim_thompson5910):

which is a permutation

OpenStudy (anonymous):

now waht

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

well i believe if i divided it be 11880/24=495 or just the permutation answer i dont know the rules so i cant really decide

jimthompson5910 (jim_thompson5910):

you have the correct answer, it's 495

jimthompson5910 (jim_thompson5910):

rule: if order matters, then use a permutation. If order does NOT matter, then use a combination

OpenStudy (anonymous):

yes ur right my answer was correct 495